The maximum current in a 26 μF capacitor connected to an ac generator with a frequency of 140 Hz is 0.15 A .

Part A

What is the maximum voltage of the generator?

Express your answer using two significant figures.

Vmax =

6.6

V ----> correct answer

Part B

What is the voltage across the capacitor when the current in the circuit is 0.10 A and increasing?

Express your answer using two significant figures.

V = ??? V

I've tried 7.4 but its wrong

Part C

What is the voltage across the capacitor when the current in the circuit is 0.10 A and decreasing?

Express your answer using two significant figures.

V = ???? V
Answer #1

a)

\omega = 2 X 3.14 X 140

\omega = 879.2 rad/sec

XC = 1 / C \omega

XC = 1 / 26 X 10-6 X 879.2

= 43.74 ohm

Imax = 0.15 A

Vmax = Imax X XC

Vmax = 0.15 X 43.74 = 6.56 V

Vmax = 6.6 volts ( rounded off value )

b )

I = Imax cos ( \omega t + \pi/2 )

I = 0.15 X sin\omega t

given current 0.1 A

0.1 = 0.15 sin\theta

\theta = 41.81o or 138.19o

V = 6.56 X cos41.81

V = 4.88 volts

c )

V = 6.63 X cos138.19

V = - 4.88 volts